OpenStudy (anonymous):

The probability of randomly selecting a red marble from a bag, replacing it, and selecting a red marble again is 1/25. What is the probability of selecting a red marble on the first try?

OpenStudy (anonymous):

\[\frac{1}{25}\]

OpenStudy (anonymous):

I will become your fan and give you a medal for the right answer

OpenStudy (anonymous):

reason: same number of marbles in the bag the second time as the first. same number of red marbles the second time as well. so nothing has changed and the probabilities are the same.

OpenStudy (anonymous):

oh thanks!

OpenStudy (anonymous):

oh wait. i need to read the question more carefully. hold on.

OpenStudy (anonymous):

hold on i think i was mistaken. sorry.

OpenStudy (anonymous):

i misread. let me try again and this time go slow, and maybe with an explanation.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

put A = the event you select a red marble on the first try. we want \[P(A)\] B = event you select a red marble on second try.

OpenStudy (anonymous):

\[A\cap B\] is the even that both marbles selected were red, and you are given that \[P(A\cap B)=\frac{1}{25}\] because you are given that the probability that both marbles selected were red

OpenStudy (anonymous):

in general \[P(A\cap B) = P(A)\times P(B|A)\] but since you have replaced the red marble there are the same number of red marble in the bag and the same number of marbles, so \[P(B)=P(A)\]

OpenStudy (anonymous):

Shouldn't it be 1/5. 1/5 * 1/5 = 1/25 th

OpenStudy (anonymous):

therefore you have \[\frac{1}{25}=P(A)\times P(A)\] so \[P(A)=\frac{1}{5}\]

OpenStudy (anonymous):

That is the probability of selecting it once into the same probability of it being picked again

OpenStudy (anonymous):

yes you are right isan_meswani. my first answer was a mistake. i read the problem incorrectly.

OpenStudy (anonymous):

1/5? not1/25?

OpenStudy (anonymous):

yes, 1/5

OpenStudy (anonymous):

sorry i messed you up on first try.

OpenStudy (anonymous):

Suppose you know there are originally 2 red marbles in the bag. How many marbles are there in all?

OpenStudy (anonymous):

Now this is simple, isn't it? its 10

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

yes that was an easy one. \[\frac{1}{5}=\frac{2}{x}\]

OpenStudy (anonymous):

oh thanks!